OpenStudy (anonymous):

You would like to have $260,000 in 18 years by making regular deposits at the end of each month in an annuity that pays an annual interest rate of 4.5% compounded monthly. How much of the $260,000 comes from interest? In your calculations, do not round until the final answer. Then, round the monthly payment to the nearest dollar.

OpenStudy (tkhunny):

i = 0.045 -- Annual Interest Rate j = i/12 -- Monthly Interest Rate r = 1+j -- Monthly Accumulation Factor n = 18*12 = 216 p = The regular payment The last payment will be ON the completion date. This payment will receive no interest. p The second to last payment will receive interest for one month, pr The third to last payment will receive interest for 2 months. Thus: \(260000 = p(1 + r + r^{2} + ... + r^{n-1}) = p\dfrac{1-r^{n}}{1-r}\) You're almost done. \(p = 260000\cdot\dfrac{1-r}{1-r^{n}}\) Amount Deposited: 216 * p Amount from Interest: 260000 - (216*p)

OpenStudy (anonymous):

Is it 783.45 ?? @tkhunny

OpenStudy (tkhunny):

Is what 783.45?

OpenStudy (tkhunny):

That looks like the right payment. Now what?